# Small Cancellation for Random Branched Covers of Groups

### v1

@@ 1. Introduction
The probabilistic method was originally pioneered by Erdös, and was used as a non-constructive approach to showing the existence of interesting examples in combinatorics and graph theory, see e.g. the classic text by Alon—Spencer [AS16]. The method has been particularly effective in graph theory, with random graphs having evolved into its own field of study. Random models have since been developed to study higher dimensional simplicial complexes (Kahle [Kah14]), random closed surfaces (Brooks—Makover [BM04]), and random $3$-manifolds (Dunfield—Thurston [DT06]). In the late 1980s, Gromov launched the study of random groups. The two main models for random groups are the density model and the few relator model, see Gromov [Gro93] and the survey article by Ollivier [Oll05].
@@
A different topological construction that is commonly used in low-dimensional topology is that of branched covers. All closed oriented surfaces can be realized as branched covers over the sphere, a fact that is also true for $3$-manifolds (Hilden [Hil74] and Montesinos [Mon74]). Branched covers have also been a source of many interesting examples in the geometry of negatively curved manifolds, see e.g. Gromov—Thurston [GT22], Fine—Premoselli [FP20], Stover—Toledo [ST22], Minemyer [Min23], Guenancia-Hamenstädt [GH25]. From this viewpoint, it is natural to look for a random model for branched covers. In the present paper, we construct a random model for branched covers of finite polygonal $2$-complexes and prove:
@@
It is easy to see that the Main Theorem fails [^1] More precisely, the theorem holds trivially when the rank is one (see Remark 3.12), and fails when two attaching maps are homotopic (these create $\pi_{2}$). The hypothesis on proper powers is necessary for the construction of our random model, see Section 3.2. for non-acceptable $2$-complexes, so our hypotheses are actually necessary.
It is well-known that the $C^{\prime}(1/6)$-small cancellation property has strong geoemetric consequences, e.g. [Gro03], [Wis04]. As a result, the $\lambda=1/6$ case of our main theorem immediately implies the following:
@@ Acknowledgments
Cho and Lafont were partially supported by the NSF, under grant DMS-2407438. Skipper was partially supported by the NSF, under grants DMS-2343739 and DMS-2506840, and the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation program (grant agreement No.725773). The authors would like to thank the following colleagues for insightful conversations and/or helpful comments: Laurent Bartholdi, Francesco Fournier-Facio, Jingyin Huang, Matthew Kahle, Ilya Kapovich, Olga Kharlampovich, MurphyKate Montee, Denis Osin, Doron Puder, Tim Riley, Genevieve Walsh, and Matthew Zaremsky.
@@ 2.1. Branched Coverings
Let us recall the notion of branched covering of smooth manifolds (see e.g. [GS23]).
@@ 2.2. Small Cancellation Conditions
Small cancellation has been a useful tool in combinatorial group theory since the 1970s, see e.g. the references [Lyn77], [Gro87], [Gro03], and [Gui12]. There are various notions of small cancellation. Here, we start by recalling the classical small cancellation with respect to a group presentation. Roughly speaking, this condition says that any common subword between two relators in a presentation is short compared to the length of the relators.
@@
[Gro87] If a finitely presented group $\Gamma$ satisfies $C^{\prime}(\frac{1}{6})$, then $\Gamma$ is word hyperbolic.
Geometric consequences of the small cancellation hypothesis are further studied in [Lyn77], typically via the group’s presentation $2$-complex, as well as van Kampen diagrams and their mapping to $2$-complexes. This allows the small cancellation condition to be reformulated geometrically, and the results to be generalized to the setting of polygonal $2$-complexes.
@@ 3.2. Random model for branched coverings
Consider the finite group $\mathbb{Z}_{2}\cong\langle x\mid x^{2}\rangle$. The presentation two complex $X$ has a single loop labelled $x$, and a single 2-cell attached by the degree two map $r$ on the circle. $X$ is homeomorphic to the projective plane $\mathbb{R}P^{2}$, a closed non-orientable surface, so any branched cover of $X$ would also have to be a closed surface. Now consider the space $X(\sigma)$, where $\sigma=(12\ldots n)$ is the cyclic permutation. The $1$-skeleton of $X(\sigma)$ is then $C_{n}$, a cycle of length $n$. There is only one lift $\tilde{r}$ of the attaching map to the $1$-skeleton $C_{n}$. Observe that the lifted map $\tilde{r}$ to $C_{n}$ is either a degree one map if $n$ is even, or a degree two map if $n$ is odd. So the space $X(\sigma)$ is either homeomorphic to $\mathbb{R}P^{2}$ (if $n$ odd) or to $\mathbb{D}^{2}$ (if $n$ even). Of these $X(\sigma)$, only the $n$ odd case produces a branched cover. In the $n$ even case, the natural map $X(\sigma)\rightarrow X$ has the property that each point has exactly $n$-preimages. The only points where it fails to be a branched cover are along the edges of $X$, where the local topology fails to be preserved.
@@
So to decide whether the canonical map $X(\sigma)\rightarrow X$ is a covering map along the interior of edges, it suffices to compare, for an edge $e\subset X^{(1)}$ and a fixed pre-image edge $e_{i}\subset X(\sigma)^{(1)}$, the number of occurrences of those edges along labels on the boundaries of the disks in $X$ and $X(\sigma)$ respectively. Let $k$ denote the number of occurrences of edge $e$, and $k_{i}$ denote the number of occurrences of the edge $e_{i}$. Our goal is to show $k_{i}=k$.
Given a $2$-cell in $X$, obtained by attaching a disk $D$, we can fix an occurrence of the edge $e$ along $\partial D$. Consider all the disks $D_{1},\ldots,D_{m}$ in $X(\sigma)$ lying above $D$. These disks are attached along all the possible lifts of the curve $\partial D$. Corresponding to these lifts, we have covering maps $\partial D_{j}\rightarrow\partial D$, and the fixed occurrence of the edge $e$ in $\partial D$ has pre-images in the $\partial D_{1},\ldots,\partial D_{m}$. Since these are all possible lifts of $\partial D$, these pre-images of the fixed occurrence of $e$ in $D$ all have distinct indices – so at most one of these lifts has label $e_{i}$. Since each occurrence of $e$ along a disk $D$ gives rise to at most one occurrence of $e_{i}$ along one of the lifts of $D$, we obtain the inequality $k_{i}\leq k$.
As a cautionary note, observe that without the hypothesis that the complex is acceptable, it is possible that none of the pre-images has label $e_{i}$ (see Example 3.6). So for non-acceptable complexes, the inequality $k_{i}\leq k$ could be strict. This is essentially due to the fact that $\sum\deg(\partial D_{i}\rightarrow\partial D)$ could potentially be smaller than $n$, so that the number of pre-images of $e$ in the $\coprod\partial D_{i}$ might be less than $n$. This also tells us that, to prove equality, we will need to make use of the “acceptable” hypothesis.
Conversely, for each distinct occurrence of $e$ along a boundary label of a disk in $X$, we can look at the attaching map for that disk, viewed as a loop starting at $e$. We can then lift that loop starting at the pre-image edge $e_{i}$, and find the attaching map for a disk in $X(\sigma)$. For a general polygonal $2$-complex, there is no guarantee that the lifted attaching maps starting at distinct occurrences of $e$ give you distinct attaching loops in $X(\sigma)$, see e.g. Example 3.6. However, any two distinct occurrences of the edge $e$ in the label of disks $D_{i},D_{j}$ form a sub-overlap. In an acceptable $2$-complex, Lemma 2.7 tells us this sub-overlap extends to a maximal sub-overlap of bounded length. In particular, the lifted loops based at the two occurrences must eventually diverge, so they correspond to distinct lifted disks in $X(\sigma)$. This implies that, in an acceptable $2$-complex, we also have the inequality $k_{i}\geq k$, which establishes the equality $k_{i}=k$.
@@
We now have, for each natural number $n$, a model that randomly produces a degree $n$ branched cover $X(\sigma)$ of the finite acceptable presentation $2$-complex $X$. We will be interested in understanding topological and geometric properties of $X(\sigma)$, as $n$ gets large.
@@ 3.3. Connectedness of branched covers
So understanding connectedness of our random branched covers is completely equivalent to understanding when a randomly selected $t$-tuple of elements in ${\text{Sym}}(n)$ generates a transitive subgroup. This is a classically studied problem, and we have the following result of Dixon [Dix69]:
@@ 3.5. Disks in random branched covering spaces
For the relators $r_{1},\dots,r_{s}$ and a random choice of permutations $\sigma=(\sigma_{1},\dots,\sigma_{t})$, $\sigma_{i}\in{\text{Sym}}(n)$, we define another type of permutation $r_{i}(\sigma)\in{\text{Sym}}(n)$ that represents the structure of lifts of the disk $D_{i}$. In $X(\sigma)$, we first attach a disk for a lift of $D_{i}$ of the relator $r_{i}$ starting from the vertex $v_{1}$. Let $v_{i(1)}:=v_{1}$ and let $v_{i^{j}(1)}$ be the vertex that is arrived at after following the letters of $r_{i}$ a total of $j$ times. Let $k\geq 1$ be the smallest such that $v_{i^{k}(1)}=v_{1}$. Then $(v_{i(1)}\cdots v_{i^{k}(1)})$ forms a cycle of length $k$ with entries corresponding to the indices of the vertices appearing in the procedure. We repeat the same process at a vertex that is not already obtained in a previous cycle until there are no vertices remaining. Finally we get a permutation of $n$ elements and denote it $r_{i}(\sigma)$. One checks the cycle lengths of $r_{i}(\sigma)$ have a one-to-one correspondence with the indices of lifts of $D_{i}$. The permutation $r_{i}(\sigma)$ is the result of applying the word map $r_{i}:{\text{Sym}}(n)\times\cdots\times{\text{Sym}}(n)\rightarrow{\text{Sym}}(n)$ (see [HP23]) to the $t$-tuple $\sigma$ of permutations. We will use the following result from [HP23, Corollary 1.5]:
@@
Let $k\geq 2$ be a fixed integer. If the permutation $r_{i}$ is not a proper power, then the expected number of cycles of length $k$ in $r_{i}(\sigma)$ is $\frac{1}{k}+\mathbf{O}(n^{-\pi(r_{i})})$, where $\pi(r_{i})\geq 2$ is the primitive rank defined in [HP23].

### v2

@@ 1. Introduction
The probabilistic method was originally pioneered by Erdös, and was used as a non-constructive approach to showing the existence of interesting examples in combinatorics and graph theory, see e.g. the classic text by Alon—Spencer [alon-spencer]. The method has been particularly effective in graph theory, with random graphs having evolved into its own field of study. Random models have since been developed to study higher dimensional simplicial complexes (Kahle [Kahle2014]), random closed surfaces (Brooks—Makover [Brooks-Makover]), and random $3$-manifolds (Dunfield—Thurston [Dunfield-Thurston]). In the late 1980s, Gromov launched the study of random groups. The two main models for random groups are the density model and the few relator model, see Gromov [Gromov-asymptotic-invariants] and the survey article by Ollivier [Ollivier-2005].
@@
A different topological construction that is commonly used in low-dimensional topology is that of branched covers. All closed oriented surfaces can be realized as branched covers over the sphere, a fact that is also true for $3$-manifolds (Hilden [hilden] and Montesinos [montesinos]). Branched covers have also been a source of many interesting examples in the geometry of negatively curved manifolds, see e.g. Gromov—Thurston [gromov-thurston], Fine—Premoselli [fine-premoselli], Stover—Toledo [stover-toledo], Minemyer [Min25], Guenancia-Hamenstädt [GH25]. From this viewpoint, it is natural to look for a random model for branched covers. In the present paper, we construct a random model for branched covers of finite polygonal $2$-complexes and prove:
@@
It is easy to see that the Main Theorem fails [^1] More precisely, the theorem holds trivially when the rank is one (see Remark 3.13), and fails when two attaching maps are homotopic (these create $\pi_{2}$, see Lemma 5.1). In the presence of proper powers, a positive proportion of the branched covers produced will have non-trivial $\pi_{2}$, see Section 5.2. for non-acceptable $2$-complexes, so our hypotheses are actually necessary.
It is well-known that the $C^{\prime}(1/6)$-small cancellation property has strong geoemetric consequences, e.g. [gromov2003random], [wise-small-cancellation]. As a result, the $\lambda=1/6$ case of our main theorem immediately implies the following:
@@
It is perhaps worth emphasizing that our model is qualitatively very different from models previously considered in the literature. Let us highlight a few of the key differences:
  • in our model, the number of generators for the random groups is growing to infinity as $n$ increases. In previous models, the number of generators was always fixed.
  • in our model, the groups that are produced virtually surject to the initial group. In previous models, the groups produced are typically quotients of the initial group.
  • in our model, the key probabilistic ingredient are asymptotic properties of random elements in finite symmetric groups. These do not appear to feature in any of the previous models.
Thus while the statements of our results are similar to those known for other models, the methods and context are quite different.
@@ Acknowledgments
Cho and Lafont were partially supported by the NSF, under grant DMS-2407438. Skipper was partially supported by the NSF, under grants DMS-2343739 and DMS-2506840, and the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation program (grant agreement No.725773). The authors would like to thank the following colleagues for insightful conversations and/or helpful comments: Laurent Bartholdi, Francesco Fournier-Facio, Jingyin Huang, Matthew Kahle, Ilya Kapovich, Olga Kharlampovich, MurphyKate Montee, Denis Osin, Tim Riley, Genevieve Walsh, and Matthew Zaremsky. The authors would also like to thank Doron Puder for comments on an earlier draft that helped to unify the main construction.
@@ 2.1. Branched Coverings
Let us recall the notion of branched covering of smooth manifolds (see e.g. [gompf20234]).
@@ 2.2. Small Cancellation Conditions
Small cancellation has been a useful tool in combinatorial group theory since the 1970s, see e.g. the references [lyndon1977combinatorial], [gromov1987hyperbolic], [gromov2003random], and [guirardel2012geometric]. There are various notions of small cancellation. Here, we start by recalling the classical small cancellation with respect to a group presentation. Roughly speaking, this condition says that any common subword between two relators in a presentation is short compared to the length of the relators.
@@
[gromov1987hyperbolic] If a finitely presented group $\Gamma$ satisfies $C^{\prime}(\frac{1}{6})$, then $\Gamma$ is word hyperbolic.
Geometric consequences of the small cancellation hypothesis are further studied in [lyndon1977combinatorial], typically via the group’s presentation $2$-complex, as well as van Kampen diagrams and their mapping to $2$-complexes. This allows the small cancellation condition to be reformulated geometrically, and the results to be generalized to the setting of polygonal $2$-complexes.
@@ 3.2. Random model for branched coverings
Consider the finite group $\mathbb{Z}_{2}\cong\langle x\mid x^{2}\rangle$. The presentation two complex $X$ has a single loop labelled $x$, and a single 2-cell attached by the degree two map $r$ on the circle. $X$ is homeomorphic to the projective plane $\mathbb{R}P^{2}$, a closed non-orientable surface, so any branched cover of $X$ would also have to be a closed surface. Now consider the space $X(\sigma)$, where $\sigma=(12\ldots n)$ is the cyclic permutation. The $1$-skeleton of $X(\sigma)$ is then $C_{n}$, a cycle of length $n$. There is only one lift $\tilde{r}$ of the attaching map to the $1$-skeleton $C_{n}$. Observe that the lifted map $\tilde{r}$ to $C_{n}$ is either a degree one map if $n$ is even, or a degree two map if $n$ is odd. So the space $X(\sigma)$ is either homeomorphic to $\mathbb{R}P^{2}$ (if $n$ odd) or to $\mathbb{D}^{2}$ (if $n$ even). Of these $X(\sigma)$, only the $n$ odd case produces a branched cover. In the $n$ even case, the natural map $X(\sigma)\rightarrow X$ has the property that each point has exactly $n$-preimages. The only points where it fails to be a branched cover are along the edges of $X$, where the local topology fails to be preserved. Note, however, that this issue can easily be fixed – one only needs to change the attaching map so that we have the correct local multiplicity along the edges. This can be done by gluing multiple copies of the lifted disk. In the case where $n=2k$ is even, attaching two copies of a disk along the same attaching map resolves the issue. This gives a space which is topologically an $S^{2}$, with the two disks identified with the two hemispheres attached to the equator, a combinatorial loop of length $2k$. In terms of the covering of the equator circle, the generator of the deck group $\mathbb{Z}_{2k}$ acts by the cyclic permutation on the labels, so by a rotation on the equatorial loop. The generator also swaps the two attached disks, so that the centers of the two disks (north and south pole) are now a pair of branch points of index $k$, interchanged by the generator. This realizes $S^{2}$ as a branched cover of degree $n=2k$ over $\mathbb{R}P^{2}$, with a single branch point in $\mathbb{R}P^{2}$ having two preimages, each with ramification $k$. We will call this a branched cover with multiplicity and give more details in Remark 3.8.
Building a branched cover with multiplicity is not the only option. Indeed, one can also change the power on an attaching map. In the case where $n=2k$ is even for instance, one can instead attach a single disk using the square of the lifted map.
@@
So to decide whether the canonical map $X(\sigma)\rightarrow X$ is a covering map along the interior of edges, it suffices to compare, for an edge $e\subset X^{(1)}$ and a fixed pre-image edge $\bar{e}\subset X(\sigma)^{(1)}$, the number of occurrences of those edges along labels on the boundaries of the disks in $X$ and $X(\sigma)$ respectively. Let $k$ denote the number of occurrences of the edge $e$, and $\bar{k}$ denote the number of occurrences of the edge $\bar{e}$. Our goal is to show $\bar{k}=k$.
Given a $2$-cell in $X$, obtained by attaching a disk $D$, we can fix an occurrence of the edge $e$ along $\partial D$. Consider all the disks $D_{1},\ldots,D_{m}$ in $X(\sigma)$ lying above $D$. These disks are attached along all the possible lifts of the curve $\partial D$. Corresponding to these lifts, we have covering maps $\partial D_{j}\rightarrow\partial D$, and the fixed occurrence of the edge $e$ in $\partial D$ has pre-images in the $\partial D_{1},\ldots,\partial D_{m}$. Since these are all possible lifts of $\partial D$, these pre-images of the fixed occurrence of $e$ in $D$ all have distinct indices – so at most one of these lifts can be $\bar{e}$. Since each occurrence of $e$ along a disk $D$ gives rise to at most one occurrence of $\bar{e}$ along one of the lifts of $D$, we obtain the inequality $\bar{k}\leq k$.
As a cautionary note, observe that without the hypothesis that the complex is acceptable, it is possible that none of the pre-images has a given label $e_{i}$ (see Example 3.6). So for non-acceptable complexes, the inequality $\bar{k}\leq k$ could be strict. This is essentially due to the fact that $\sum\deg(\partial D_{i}\rightarrow\partial D)$ could potentially be smaller than $n$, so that the number of pre-images of $e$ in the $\coprod\partial D_{i}$ might be less than $n$. This also tells us that, to prove equality, we will need to make use of the “acceptable” hypothesis.
Conversely, for each distinct occurrence of $e$ along a boundary label of a disk in $X$, we can look at the attaching map for that disk, viewed as a loop starting at $e$. We can then lift that loop starting at the chosen pre-image edge $\bar{e}$, and find the attaching map for a disk in $X(\sigma)$. For a general polygonal $2$-complex, there is no guarantee that the lifted attaching maps starting at distinct occurrences of $e$ give you distinct attaching loops in $X(\sigma)$, see e.g. Example 3.6. However, any two distinct occurrences of the edge $e$ in the label of disks $D_{i},D_{j}$ form a sub-overlap. In an acceptable $2$-complex, Lemma 2.7 tells us this sub-overlap extends to a maximal sub-overlap of bounded length. In particular, the lifted loops based at the two occurrences must eventually diverge, so they correspond to distinct lifted disks in $X(\sigma)$. This implies that, in an acceptable $2$-complex, we also have the inequality $\bar{k}\geq k$, which establishes the equality $\bar{k}=k$.
@@
Remark 3.8.
We conclude this section by discussing how to extend our random model to the case where the original $2$-complex $X$ is not acceptable. Consider the situation where a disk is attached along a map that is a proper power in $\pi_{1}(X^{(1)})$, e.g. the attaching map is along $\alpha=w^{k}$ for some word $w$ and $k\geq 2$. We may assume that $w$ itself is non-trivial, and not a proper power. As was discussed earlier, the key problem that can arise are issues with the multiplicity of the lifted disk along edges. To analyze this, consider a lift $\tilde{w}$ of the loop $w$ of degree $d$. Then the corresponding lift $\tilde{\alpha}$ of the attaching map is the power $\tilde{w}^{r}$, where $r=LCM(k,d)$. As a covering of $\alpha$, this lifted map has degree $r/k$ – which can potentially be different from the desired degree $d$. As discussed in Example 3.6, there are several ways to resolve the local multiplicity issue, but we will focus on just one: attaching multiple copies of the lifted disk, each along the map $\tilde{\alpha}$. In order to get degree $d$, we need to attach $d/(r/k)=kd/r=dk/LCM(k,d)=GCD(k,d)$ copies of the disk along the same map $\tilde{\alpha}$. We call this the random model with multiplicity.
We now have, for each natural number $n$, models that randomly produces a degree $n$ branched cover $X(\sigma)$ of the finite presentation $2$-complex $X$. We will be interested in understanding topological and geometric properties of $X(\sigma)$, as $n$ gets large.
@@ 3.3. Connectedness of branched covers
So understanding connectedness of our random branched covers is completely equivalent to understanding when a randomly selected $t$-tuple of elements in ${\text{Sym}}(n)$ generates a transitive subgroup. This is a classically studied problem, and we have the following result of Dixon [dixon1969probability]:
@@ 3.5. Disks in random branched covering spaces
For the relators $r_{1},\dots,r_{s}$ and a random choice of permutations $\sigma=(\sigma_{1},\dots,\sigma_{t})$, $\sigma_{i}\in{\text{Sym}}(n)$, we define another type of permutation $r_{i}(\sigma)\in{\text{Sym}}(n)$ that represents the structure of lifts of the disk $D_{i}$. In $X(\sigma)$, we first attach a disk for a lift of $D_{i}$ of the relator $r_{i}$ starting from the vertex $v_{1}$. Let $v_{i(1)}:=v_{1}$ and let $v_{i^{j}(1)}$ be the vertex that is arrived at after following the letters of $r_{i}$ a total of $j$ times. Let $k\geq 1$ be the smallest such that $v_{i^{k}(1)}=v_{1}$. Then $(v_{i(1)}\cdots v_{i^{k}(1)})$ forms a cycle of length $k$ with entries corresponding to the indices of the vertices appearing in the procedure. We repeat the same process at a vertex that is not already obtained in a previous cycle until there are no vertices remaining. Finally we get a permutation of $n$ elements and denote it $r_{i}(\sigma)$. One checks the cycle lengths of $r_{i}(\sigma)$ have a one-to-one correspondence with the indices of lifts of $D_{i}$. The permutation $r_{i}(\sigma)$ is the result of applying the word map $r_{i}:{\text{Sym}}(n)\times\cdots\times{\text{Sym}}(n)\rightarrow{\text{Sym}}(n)$ (see [hanany2023word]) to the $t$-tuple $\sigma$ of permutations. We will use the following result from [hanany2023word, Corollary 1.5]:
@@
Let $k\geq 2$ be a fixed integer. If the permutation $r_{i}$ is not a proper power, then the expected number of cycles of length $k$ in $r_{i}(\sigma)$ is $\frac{1}{k}+\mathbf{O}(n^{-\pi(r_{i})})$, where $\pi(r_{i})\geq 2$ is the primitive rank defined in [hanany2023word].
@@ 5.1. Multiple vertex case
5.2. Non-acceptable $2$-complexes
In Section 3.2, we discussed a possible extension of our random model to general $2$-complexes (not necessarily acceptable). To understand the topology of a random branched cover in the non-acceptable case, the following observation will be useful.
Lemma 5.1.
Let $X$ be a $2$-complex, where a pair of $2$-cells are attached along homotopic maps $\alpha\simeq\beta:S^{1}\rightarrow X^{(1)}$. Then $\pi_{2}(X)\neq 0$, so $X$ is not aspherical.
Proof.
Up to homotopy equivalence, we can assume that $\alpha=\beta$, and that these are the last two disks attached. Denote by $X_{\beta}$ the space obtained from $X$ by removing the disk attached along $\beta$. Then the attaching map $\beta$, viewed as a map into $X_{\alpha}$, can be homotoped across the disk attached along $\alpha$ to a point map. It follows that $X$ is homotopy equivalent to the join $X_{\beta}\vee S^{2}$. Applying the $\pi_{2}$-functor, and noting that $S^{2}$ is a retract of $X_{\beta}\vee S^{2}$, we see that
$\mathbb{Z}\cong\pi_{2}(S^{2})\hookrightarrow\pi_{2}(X_{\beta}\vee S^{2})\cong\pi_{2}(X)$
establishing the result. ∎
For the random model with multiplicity, any disk whose attaching map is a proper power $\alpha^{k}$ ($\alpha$ not a proper power, and $k\geq 2$) can give rise to multiple disks in $X(\sigma)$ attached along the same lifted map $\tilde{\alpha}$. This occurs anytime the degree $d$ of the lifted map $\tilde{\alpha}\rightarrow\alpha$ has a divisor in common with the power $k$. In those situations, Lemma 5.1 applies and guarantees that $X(\sigma)$ is not aspherical. Note that the degrees $d$ of the lifted map are the lengths of the cycles in the image of the word map $\alpha(\sigma)$. Since $\alpha$ is not a proper power, the permutations $\alpha(\sigma)$ are uniformly distributed in the symmetric group as $\sigma$ ranges over all tuples of permutations. In particular, for $n$ sufficiently large, there is a positive proportion of $\sigma$ that will produce a permutation $\alpha(\sigma)$ containing a $k$-cycle. This gives us the following
Corollary 5.2.
Assume the $2$-complex $X$ has a disk where the attaching map is a proper power $\alpha^{k}$ in $\pi_{1}(X^{(1)})$. Then for $n$ large, a positive proportion of the random branched covers $X(\sigma)$ in the random model with multiplicity will satisfy $\pi_{2}(X(\sigma))\neq 0$.